# V-FLUTE: Visual Figurative Language Understanding with Textual Explanations

### v1

@@ Abstract
Large Vision-Language models (VLMs) have demonstrated strong reasoning capabilities in tasks requiring a fine-grained understanding of literal images and text, such as visual question-answering or visual entailment. However, there has been little exploration of these models’ capabilities when presented with images and captions containing figurative phenomena such as metaphors or humor, the meaning of which is often implicit. To close this gap, we propose a new task and a high-quality dataset: Visual Figurative Language Understanding with Textual Explanations (V-FLUTE). We frame the visual figurative language understanding problem as an explainable visual entailment task, where the model has to predict whether the image (premise) entails a claim (hypothesis) and justify the predicted label with a textual explanation. Using a human-AI collaboration framework, we build a high-quality dataset, V-FLUTE, that contains 6,027 image, claim, label, explanation instances spanning five diverse multimodal figurative phenomena: metaphors, similes, idioms, sarcasm, and humor. The figurative phenomena can be present either in the image, the caption, or both. We further conduct both automatic and human evaluations to assess current VLMs’ capabilities in understanding figurative phenomena.111Code and data will be available at github.com/asaakyan/V-FLUTE
@@ 1 Introduction
Figurative language is integral to human communication, enabling a variety of communicative goals Roberts and Kreuz (1994), including affective communication Fussell and Moss (2014). Figurative language presents a significant challenge to computational approaches as it requires understanding of implicit meaning behind an expression Stowe et al. (2022); Shutova (2011); Veale et al. (2016); Zhou et al. (2021). Recently, Chakrabarty et al. (2022) proposed a task and dataset for Figurative Language Understanding through Textual Explanations (FLUTE) that frames the problem as an explainable textual entailment covering a variety of figurative language phenomena in text: metaphors, similies, idioms, and sarcasm. This dataset has been used successfully to advance and benchmark the capabilities of LLMs for understanding figurative language in text Saakyan et al. (2022); Ziems et al. (2024); Sravanthi et al. (2024); Dey et al. (2024).
However, figurative meaning is also prevalent in visual phenomena, such as visual metaphors Akula et al. (2023); Chakrabarty et al. (2023), multimodal sarcasm Desai et al. (2022), and humor Hessel et al. (2023); Hwang and Shwartz (2023). Yet so far most of the work on vision and language models (VLMs) has focused on understanding literal meaning in images and captions (e.g., ScienceQA Lu et al. (2022), MMMU Yue et al. (2024)) including work on explainable visual entailment Kayser et al. (2021). Building on the idea of FLUTE Chakrabarty et al. (2022) for text, we present a new dataset for visual figurative language understanding with textual explanations (V-FLUTE). Our dataset contains 6,027 image, claim, label, explanation instances spanning diverse figurative phenomena. Each instance contains an image (premise) and a textual claim (hypothesis) that is either entailed or contradicted by the image. Deciding the entailment relation requires the vision-language model to understand the implicit meaning in both the visual and textual modalities. Our dataset contains figurative phenomena present in the image, in the caption, or in both. In addition, to mitigate the dependence on spurious correlations, to more rigorously investigate reasoning capabilities, and to promote explainability, our task requires the model to generate a plausible explanation for the output label. See Figure 1 for two examples from our dataset.
We make the following contribution towards assessing VLMs ability to understand multimodal figurative phenomena:
V-FLUTE, a high-quality dataset of 6,027 image, claim, label, explanation instances built using a human-LLMs collaboration framework covering several phenomena: metaphors, similies, idioms, sarcasm, and humor (Section 3). We will make the dataset available.
A detailed human evaluation with error analysis yielding insights into types of errors for different classes of models (Section 5).
@@ 2 Related Work
Phenomenon | Data Source | Figurative Part | Our Contribution | # instances | |||||
@@
Image+Caption | Taken As Is | 651 (651 E) |
Textual entailment MacCartney and Manning (2008); Bowman et al. (2015) and visual entailment Xie et al. (2019) tasks have been proposed to measure language and multimodal understanding. However, models trained to simply improve label accuracy on these data can be brittle and suffer from spurious correlations Poliak et al. (2018); Gururangan et al. (2018); McCoy et al. (2019); Gardner et al. (2021). Datasets such as e-SNLI Camburu et al. (2018) and e-SNLI-VE Kayser et al. (2021) augment existing entailment datasets with natural language explanations and train models to not only predict the label, but also generate a textual explanation for the reason behind the prediction. Such approach has been further adopted for a variety of tasks, such as commonsense reasoning Rajani et al. (2019); Aggarwal et al. (2021) and social norm understanding CH-Wang et al. (2023) among others Wiegreffe and Marasovic (2021). This approach has been extended to assess LLMs’ capabilities on understanding figurative language through the FLUTE dataset Chakrabarty et al. (2022). FLUTE frames figurative language understanding as an explainable textual entailment task. Recent progress in multimodal models Li et al. (2022); Alayrac et al. (2022); OpenAI (2023); Team et al. (2023); Liu et al. (2023b); Anthropic (2024) prompts us to asses similar capabilities when extended to multimodal setting, testing the understanding of non-literal meaning contained in both images and text. We present an equivalent of the FLUTE dataset for the visual modality: V-FLUTE.
A number of previous works has focused on modeling figurative phenomena beyond text. Chakrabarty et al. (2023) use a human-AI collaboration framework to generate visual metaphors from linguistic metaphors (HAIVMet dataset) and propose a visual entailment task as an extrinsic evaluation of dataset quality. The dataset contains images, claims, and labels, but no textual explanations. Yosef et al. (2023) proposed a benchmark (IRFL) where given an idiom, metaphor, or simile the model has to distinguish which of the four associated images implies the figurative meaning of the expression. This dataset focuses on the figurative meaning in the textual modality and does not contain textual explanations. There has also been work on understanding multimodal sarcasm with explanations Desai et al. (2022), mostly containing noisy user-generated text and crowdworker-written explanations. Other line of work has focused on understanding humor with multimodal models. MemeCap Hwang and Shwartz (2023) is a dataset for understanding memes. Hessel et al. (2023) release a corpus of annotated New Yorker Caption Contest entries, where the goal is to come up with a humorous captions for an image, with high-quality explanations for why the caption is humorous. The dataset is relatively limited in size containing only 520 unique instances in its training set. We leverage all these benchmarks to build V-FLUTE.
HAIVMet | IRFL | MuSE | MemeCap | NYCartoons | |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
Contradiction | Entailment | Contradiction | Entailment | Entailment | |||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
@@ 3 V-FLUTE Task and Dataset
Following prior work on figurative language understanding in text defined as explainable textual entailment Chakrabarty et al. (2022), we define the visual figurative understanding as an explainable visual entailment task: given an image (premise) and a claim (hypothesis) , output a textual explanation justifying whether the premise entails or contradicts the hypothesis and assign a label . We focus on the binary classification task since, for neutral labels, the explanations would be trivial (simply describing the image).
To build V-FLUTE, we start with existing multimodal figurative datasets and use human-AI collaboration frameworks with expert annotators Chakrabarty et al. (2022); Wiegreffe et al. (2022); Liu et al. (2022) to transform them into a high-quality, explainable visual entailment benchmark. These datasets cover particular phenomena such as metaphors, similes, idioms, sarcasm or humor. Each instance includes an image and a caption and the figurative phenomenon can be either in the image, the caption or in both. We transform each data into a unified image, claim, label, explanation format for explainable visual entailment. An overview of the dataset and our contributions can be found in Table 1. See examples from each dataset in Table 2. Below, we describe the construction of V-FLUTE for each figurative language type (metaphors & similes, idioms, sarcasm and humor).
3.1 Metaphors and Similes
Metaphors and similes are powerful rhetorical devices that can be expressed either in text or visually in an image. Visual metaphors are used as persuasive devices in various fields such as advertising Forceville (2002); Scott (1994). To create visual entailment instances containing metaphors and similes in V-FLUTE, we rely on two existing resources: HAIVMet Chakrabarty et al. (2023) and IRFL Yosef et al. (2023). Instances taken from HAIVMet contain the metaphor/simile as a part of the premise (image), while those taken from IRFL have the metaphor/simile as a part of the hypothesis (text).
@@ 3.1.1 HAIVMet as Data Source
The HAIVMet Chakrabarty et al. (2023) data consists of 1,193 images of visual metaphors spanning over 958 distinct linguistic metaphors. Each image is associated with a claim that can be contradicting or entailing the image. In addition, each image is associate with a visual elaboration that presents a textual description of the image (See Figure 2). This visual elaboration was used in the original paper to generate the visual metaphors (images).
Generating Textual Explanations. We augment the dataset with candidate textual explanations. We prompt ChatGPT (gpt-3.5-0914) to generate an explanation for every tuple visual elaboration, claim, label (See Figure 2; and prompt in Appendix D.1.1).
Expert Verification. Each claim is paired with up to images. However, since these images were automatically generated with DALLE-2 using the visual elaborations, not all are completely faithful. Moreover, some claims and labels were inconsistent. Finally, automatically generated LLM candidate explanations are not always correct and require refining. To tackle these issues, we employ an expert verification process involving three expert annotators with significant experience in figurative language and visual metaphor understanding. Since each claim can be paired with more than one visual metaphor, we ask annotators to select the visual metaphor most faithful to the linguistic metaphor and visual elaboration (see Image Selection in Figure 2) or select none in the rare case when none of the visual metaphors are of good quality. As a part of the same annotation round, we also ask them to verify and edit the explanation if necessary to ensure correctness and high quality. Post strict quality control, we have 857 image, claim, label, explanation instances.
3.1.2 IRFL as Data Source
The IRFL dataset Yosef et al. (2023) contains 1,440 figurative expressions, each associated with distinct images. One of those images represents the figurative expression (see Figure 3), and the other 3 act as distractors.
Generating Textual Explanations. We prompt GPT-4 (gpt-4-vision-preview) with the ground truth label, claim, and the image to explain the relationship between the image and the claim.
Expert Verification. We recruit the same three expert annotators from HAIVMET annotations and ask them to verify the explanation is adequate and edit it when necessary. We also ask the annotator to discard rare noisy instances where the claim, image, and label do not fit. Post strict quality control, we are left with 1149 image, claim, label, explanation instances.
3.2 Idioms
The IRFL dataset contains idioms in addition to metaphors and similies. An identical procedure to the one described in Section 3.1.2 was used for generating V-FLUTE instances for idioms (370 image, claim, label, explanation examples).
@@ 3.3 Sarcasm
To create visual entailment instances containing sarcasm, we rely on the MuSE data Desai et al. (2022). Similarly to IRFL, instances from MuSE data contain sarcasm in the hypothesis (text).
@@ 3.3.1 MuSE as Data Source
The MuSE dataset Desai et al. (2022) consists of 3510 distinct images, the respective sarcastic claims that act as contradiction instances (see example in Figure 4), and crowd worker written explanations justifying the contradiction.
Generating Entailment Claims. Since the dataset only contains sarcastic instances, there are no claims with an entailment relationship. We generate the entailing claims by prompting GPT-4 to generate a non-sarcastic version of the claim while maintaining the user-generated informal style of the text (see the generated entailment claim in Figure 4).
Generating Textual Explanations. While the dataset already contains crowdworker-written explanations, upon inspection, they were often deemed poor quality, lacking enough details, and formulaic (e.g., see the crowdworker explanation in Figure 4). To improve their quality, we use the dataset’s existing crowdworker explanations and prompt GPT-4 to rewrite and generate high-quality candidate textual explanations given the claim and the label (see the re-written explanation in Figure 4). See the prompt in Appendix D.3.
Expert Verification. Each image is now paired with a GPT-4-generated entailing claim, an original contradicting claim, and their respective labels and explanations. The same three expert annotators checked if the generated explanations are adequate (i.e., complete, correct, and concise) and if not, edit them. Experts were also instructed to discard noisy examples, e.g. when the image does not contradict the sarcastic claim. Through strict quality control, we obtain 1,042 image, claim, label, explanation instances.
@@ 3.4.1 MemeCap as Data Source
Claim Generation. Since meme captions are not suited for an entailment task, we perform prompt GPT-4 with the caption to generate a claim from it (see example in Figure 5). We filter these set of samples further with GPT-4 by asking whether the image entails the claim and only selecting positive instances. In addition to generating claims that entail the meme, we generate counterclaims using GPT-4.
Generating Textual Explanations. We prompted GPT-4 with the ground truth label in the prompt to explain the relationship between the image and the claim. See prompts in Appendix D.4.
Expert Verification. We hire the same three expert annotators to ensure the correctness of the data. Each annotator is tasked with verifying that 1) the generated claim fits the image and 2) the explanation is correct and complete, and if not, make the necessary changes. We also ask to discard samples with inappropriate content. After careful quality control, we have 1958 image, claim, label, explanation instances.
@@ 3.4.2 NYCartoons as Data Source
The NYCartoons dataset Hessel et al. (2023) contains 651 high-quality instances from the New Yorker Cartoon Caption Contest. Each instance consists of a humorous image paired with a caption and a natural language explanation justifying the implicit humor between the caption and the image. We simply use the existing data where the caption is treated as a claim entailing the humorous image paired with an explanation.
@@ 3.5 Dataset Statistics
We split our data into 4,578 training, 726 validation, and 723 testing instances. Detailed counts per phenomenon and dataset, as well as other statistics, are in Appendix A.

### v2

# Understanding Figurative Meaning through Explainable Visual Entailment

@@ Abstract
Large Vision-Language Models (VLMs) have demonstrated strong capabilities in tasks requiring a fine-grained understanding of literal meaning in images and text, such as visual question-answering or visual entailment. However, there has been little exploration of these models’ capabilities when presented with images and captions containing figurative meaning, such as metaphors or humor. To close this gap, we propose a new task framing the figurative meaning understanding problem as an explainable visual entailment task, where the model has to predict whether the image (premise) entails a caption (hypothesis) and justify the predicted label with a textual explanation. The figurative phenomena can be present either in the image, the caption, or both. Utilizing a human-AI collaboration approach, we build the accompanying expert-verified dataset V-FLUTE, containing 6,027 image, caption, label, explanation instances spanning five diverse figurative phenomena: metaphors, similes, idioms, sarcasm, and humor. Through automatic evaluation, we find that VLMs struggle to generalize from literal to figurative meaning, particularly when it is present in images. Further, we identify common types of errors in VLM reasoning via human evaluation.111Code and data: github.com/asaakyan/V-FLUTE
Understanding Figurative Meaning through Explainable Visual Entailment
Arkadiy Saakyan1 Shreyas Kulkarni1 Tuhin Chakrabarty1 Smaranda Muresan1 1Columbia University a.saakyan@cs.columbia.edu
@@ 1 Introduction
Figurative language is integral to human communication, enabling a variety of communicative goals Roberts and Kreuz (1994), including affective communication Fussell and Moss (2014). Figurative language presents a significant challenge to computational approaches as it requires understanding of implicit meaning behind an expression Stowe et al. (2022); Shutova (2011); Veale et al. (2016); Zhou et al. (2021). Recently, Chakrabarty et al. (2022) proposed a task and dataset for Figurative Language Understanding through Textual Explanations (FLUTE) that frames the problem as an explainable textual entailment covering a variety of figurative language phenomena in text: metaphors, similes, idioms, and sarcasm. This dataset has been used successfully to advance and benchmark the capabilities of LLMs for understanding figurative language in text Saakyan et al. (2022); Ziems et al. (2024); Sravanthi et al. (2024); Dey et al. (2024).
However, figurative meaning is also prevalent in visual phenomena, such as visual metaphors Akula et al. (2023); Chakrabarty et al. (2023), multimodal sarcasm Desai et al. (2022), and humor Hessel et al. (2023); Hwang and Shwartz (2023). Yet so far most of the work on vision and language models (VLMs) has focused on understanding literal meaning in images and captions (e.g., ScienceQA Lu et al. (2022), MMMU Yue et al. (2024)) including work on explainable visual entailment Kayser et al. (2021). Building on the idea of FLUTE Chakrabarty et al. (2022) for text, we present a new dataset for understanding figurative meaning as explainable visual entailment, we call V-FLUTE. Our dataset contains 6,027 image, caption, label, explanation instances spanning diverse figurative phenomena. Each instance contains an image (premise) and a caption (hypothesis) that is either entailed or contradicted by the image. Deciding the entailment relation requires the vision-language model to understand the implicit meaning in both the visual and textual modalities. Our dataset contains figurative phenomena present in the image, in the caption, or in both. In addition, to mitigate the dependence on spurious correlations, to more rigorously investigate reasoning capabilities, and to promote explainability, our task requires the model to generate a plausible explanation for the output label. See Figure 1 for two examples from our dataset.
We make the following contributions towards assessing VLMs ability to understand figurative meaning expressed multimodally:
V-FLUTE, an expert-verified dataset of 6,027 image, caption, label, explanation instances built using a human-LLM collaboration framework covering several phenomena: metaphors, similes, idioms, sarcasm, and humor (Section 3).
A detailed human evaluation with error analysis yielding insights into the types of errors for different classes of models (Section 5).
@@ 2 Related Work
| Phenomenon | Data Source | Visual Style | Figurative Part | Our Contribution | # instances | |||||
@@
Illustration | Image+Caption | Taken As Is | 651 (651 E) |
Textual entailment MacCartney and Manning (2008); Bowman et al. (2015) and visual entailment Xie et al. (2019) tasks have been proposed to measure language and multimodal understanding. However, models trained to simply improve label accuracy on these data can be brittle and suffer from spurious correlations Poliak et al. (2018); Gururangan et al. (2018); McCoy et al. (2019); Gardner et al. (2021). Datasets such as e-SNLI Camburu et al. (2018) and e-SNLI-VE Kayser et al. (2021) augment existing entailment datasets with natural language explanations and train models to not only predict the label, but also generate a textual explanation for the reason behind the prediction. However, they only focus on literal meaning in text and images. Recently, explainable entailment has been utilized to assess LLMs’ capabilities on understanding figurative language through the FLUTE dataset Chakrabarty et al. (2022). FLUTE frames figurative language understanding as an explainable textual entailment task. Recent progress in multimodal models Li et al. (2022); Alayrac et al. (2022); OpenAI (2023); Team (2023); Liu et al. (2023b); Anthropic (2024) prompts us to asses understanding of figurative meaning present in the multimodal setting, contained in both images and text beyond intent and sentiment Zhang et al. (2021); Kruk et al. (2019). To this end, we present an equivalent of the FLUTE dataset for the visual modality: V-FLUTE.
| HAIVMet | IRFL | MuSE | MemeCap | NYCartoons | |||||||||||||||||||||||||||||||||||||||||||
| Contradiction | Entailment | Contradiction | Entailment | Entailment | |||||||||||||||||||||||||||||||||||||||||||
@@ 3 V-FLUTE Task and Dataset
Following prior work on figurative language understanding in text defined as explainable textual entailment, FLUTE Chakrabarty et al. (2022), we define understanding figurative meaning as an explainable visual entailment task: given an image (premise) and a caption (hypothesis) , output a textual explanation justifying whether the premise entails or contradicts the hypothesis and assign a label . We focus on the binary classification task, since for neutral labels, the explanations would be trivial (simply describing the image).
To build V-FLUTE, we start with existing multimodal figurative datasets which cover phenomena such as metaphors, similes, idioms, sarcasm or humor. We utilize human-AI collaboration frameworks with expert annotators Chakrabarty et al. (2022); Wiegreffe et al. (2022); Liu et al. (2022) to augment them with expert-verified textual explanations and entailing/contradicting captions. Each instance then includes an image and a caption, and the figurative phenomenon can be either in the image, the caption or in both. An overview of V-FLUTE dataset and our contributions w.r.t to the source datasets can be found in Table 1. See examples corresponding to each source dataset in Table 2 as they appear in V-FLUTE. Below, we describe the construction of V-FLUTE by each phenomenon.
3.1 Metaphors, Similes and Idioms
To create visual entailment instances containing metaphors and similes in V-FLUTE, we rely on two existing resources: HAIVMet Chakrabarty et al. (2023) and IRFL Yosef et al. (2023). Instances from HAIVMet contain the metaphor/simile as a part of the premise (image), while those taken from IRFL have the metaphor/simile as a part of the hypothesis (text).
3.1.1 IRFL as Data Source
Yosef et al. (2023) proposed a benchmark (IRFL) where given a metaphor, a simile or an idiom the model has to distinguish which of the four associated images implies the figurative meaning of the expression. This dataset contains 1,440 figurative expressions, each associated with distinct images. One of those images represents the figurative expression (see Figure 2), and the other 3 act as distractors.
Generating Textual Explanations. We prompt GPT-4 (gpt-4-vision-preview) with the ground truth label, caption, and the image to explain the relationship between the image and the caption.
Expert Verification. We recruit three expert annotators with significant experience in figurative language and visual metaphor understanding on Upwork and ask them to verify the explanation is correct, complete, and concise and if not, edit it (see details in Appendix B). We also ask the annotators to discard rare noisy instances where the caption, image, and label do not fit (due to automatic image selection). Due to relative simplicity of generating the explanation given a literal image, the experts only needed to edit % of the explanations. They also removed the data, resulting in 1149 image, caption, label, explanation instances for metaphors and similes and 370 for idioms.
@@ 3.1.2 HAIVMet as Data Source
Chakrabarty et al. (2023) use a human-AI collaboration framework to generate visual metaphors from linguistic metaphors (HAIVMet dataset) and propose a visual entailment task as an extrinsic evaluation of dataset quality. The HAIVMet data consists of 1,193 images of visual metaphors spanning over 958 distinct linguistic metaphors. Each image is associated with a caption that can be contradicting or entailing the image. In addition, each image is associate with a visual elaboration that presents a textual description of the image (See Appendix A, Figure 7). This visual elaboration was used in the original paper to generate the visual metaphors (images).
Generating Textual Explanations. We augment the dataset with candidate textual explanations. We prompt ChatGPT (gpt-3.5-0914) to generate an explanation for every tuple visual elaboration, caption, label (See Appendix A, Figure 7; and prompt in Appendix F.1.1).
Expert Verification. Each caption is paired with up to images. However, since these images were automatically generated with DALLE-2 using the visual elaborations, not all are completely faithful. Moreover, some captions and labels were inconsistent. Finally, automatically generated LLM candidate explanations are not always correct and require refining. To tackle these issues, we employ an expert verification process recruiting the same three expert annotators as from the IRFL section above (see details in Appendix B). We ask the annotators to select the visual metaphor most faithful to the linguistic metaphor and the visual elaboration (see Image Selection in Appendix A, Figure 7) or if none were. In addition, we ask them to verify and edit the explanation if necessary to ensure correctness, completeness, and conciseness. On average, experts edited of the explanations and of captions, and rejected of visual metaphors, resulting in 857 image, caption, label, explanation instances.
@@ 3.2 Sarcasm
To create visual entailment instances containing sarcasm, we rely on the MuSE data Desai et al. (2022).
@@ 3.2.1 MuSE as Data Source
The MuSE dataset Desai et al. (2022) consists of 3510 distinct images, the respective sarcastic captions that act as contradiction instances (see example in Figure 3), and crowd worker written explanations justifying the contradiction.
Generating Entailment Captions. Since the dataset only contains sarcastic instances, there are no captions with an entailment relationship. We generate the entailing captions by prompting GPT-4 to generate a non-sarcastic version of the caption while maintaining the user-generated informal style of the text (see the generated entailment caption in Figure 3).
Generating Textual Explanations. While the dataset already contains crowdworker-written explanations, upon inspection, they were often deemed poor quality, lacking enough details, and formulaic (e.g., see the crowdworker explanation in Figure 3). To improve their quality, we use the dataset’s existing crowdworker explanations and prompt GPT-4 to rewrite and generate candidate textual explanations given the caption and the label (see the re-written explanation in Figure 3). See the prompt in Appendix F.3.
Expert Verification. Each image is now paired with a GPT-4-generated entailing caption, an original contradicting caption, and their respective labels and explanations. The same three expert annotators checked if the generated explanations are adequate (i.e., complete, correct, and concise) and if not, asked to edit them. The experts were also instructed to discard noisy examples, e.g. when the image does not contradict the sarcastic caption. On average, experts edited of the initial explanations and rejected of the examples, resulting in 1,042 image, caption, label, explanation instances.
@@ 3.3.1 MemeCap as Data Source
Caption Generation. Meme captions are not suited for an entailment task, so we prompt GPT-4 with the original caption to generate an entailing caption in the form of a claim from it (see example in Figure 4). We filter these set of samples further with GPT-4 by asking whether the image entails the caption and only selecting positive instances. In addition to generating captions that entail the meme, we generate contradicting captions using GPT-4.
Generating Textual Explanations. We prompted GPT-4 with the ground truth label in the prompt to explain the relationship between the image and the caption. See prompts in Appendix F.4.
Expert Verification. We hire the same three expert annotators to ensure the correctness of the data. Each annotator is tasked with verifying that 1) the generated caption fits the image and 2) the explanation is correct and complete, and if not, make the necessary changes. We also ask to discard samples with inappropriate content. Experts edited of the explanations and of captions on average, and discarded of inappropriate instances, resulting in 1958 image, caption, label, explanation instances.
@@ 3.3.2 NYCartoons as Data Source
The NYCartoons dataset Hessel et al. (2023) contains 651 high-quality instances from the New Yorker Cartoon Caption Contest. Each instance consists of an image paired with a humorous caption and an explanation of why this combination of the caption and the image is funny. We utilize this data as is by treating the image as entailing the caption, so the explanation of the entailment relationship is the explanation of the joke.
@@ 3.4 Dataset Statistics
We split our data into 4,578 training, 726 validation, and 723 testing instances. Detailed counts per phenomenon and dataset, as well as other statistics, are in Appendix C.